A very very thin charged shell is set into spin (at some rate linearly increasing with time ). Calculate the electric field inside ? 


1
Expert's answer
2021-01-27T09:30:59-0500

The electric field inside a shell of radius "R" and surface charge "\sigma" at distance "x" from the centre is


"E=\frac{1}{4\pi\epsilon_0}\frac{\sigma\pi R^2}{x^2}=\frac{\sigma R^2}{4\epsilon_0x^2}."


The field is 0 at the centre.